In how many ways can a committee of nine be divided into two subcommittee, one having five members and the other having four?


1
Expert's answer
2021-07-27T17:08:46-0400
"5+4=9"

"\dbinom{9}{5}=\dfrac{9!}{5!(9-5)!}=126"

"\dbinom{9}{4}=\dfrac{9!}{4!(9-4)!}=126"

A committee of nine can be divided into two subcommittee, one having five members and the other having four in 126 ways.